OpenStudy (anonymous):

The probability of an event occurring is 0.35. what are the odds in favor of the event? I think I have the answer but can someone tell me what they came up with so I can see if I got the same answer

OpenStudy (anonymous):

\[ \frac{0.35}{1-0.35} = \frac {0.35}{ 0.65} = \frac{7}{13} \]

Directrix (directrix):

@swetzbar --> I'm sticking with the answer I gave you the first time: 35: 65 which does simplify to 7:13 as shown by FFM above.